# The Sparse Frontier: Sparse Attention Trade-offs in Transformer LLMs

### v1

@@ Abstract
Sparse attention offers a promising strategy to extend long-context capabilities in Transformer LLMs, yet its viability, its efficiency–accuracy trade-offs, and systematic scaling studies remain unexplored. To address this gap, we perform a careful comparison of training-free sparse attention methods at varying model scales, sequence lengths, and sparsity levels on a diverse collection of long-sequence tasks —including novel ones that rely on natural language while remaining controllable and easy to evaluate. Based on our experiments, we report a series of key findings: 1) an isoFLOPS analysis reveals that for very long sequences, larger and highly sparse models are preferable to smaller and dense ones. 2) The level of sparsity attainable while statistically guaranteeing accuracy preservation is higher during decoding than prefilling, and correlates with model size in the former. 3) There is no clear strategy that performs best across tasks and phases, with different units of sparsification or budget adaptivity needed for different scenarios. Even moderate sparsity levels often result in significant performance degradation on at least one task, highlighting that sparse attention is not a universal solution. 4) We introduce and validate novel scaling laws specifically tailored for sparse attention, providing evidence that our findings are likely to hold true beyond our range of experiments. Through these insights, we demonstrate that sparse attention is a key tool to enhance the capabilities of Transformer LLMs for processing longer sequences, but requires careful evaluation of trade-offs for performance-sensitive applications.
$\star$\$$\star$\$footnotetext: Research conducted during an internship at Cohere. Correspondence email: piotr.nawrot@ed.ac.uk00footnotetext: Work done prior to joining Meta.
@@ 1 Introduction
The ability to model long sequences in large language models (LLMs) lies at the heart of long-context processing (Liu et al., 2025a) and inference-time scaling (Snell et al., 2024; Muennighoff et al., 2025). The fundamental bottleneck for this ability is the self-attention mechanism (Bahdanau et al., 2015) in Transformer (Vaswani et al., 2017) LLMs: during inference-time prefilling, the complexity of attention scales quadratically with sequence length—hence, ballooning time-to-first-token and deployment cost (Jiang et al., 2024). During inference-time decoding, dense attention results in a linearly growing cache of key–value (KV) items, whose size is proportional to the sequence length. Hence, runtime is dominated by high bandwidth memory access for loading these KV pairs from memory (Nawrot et al., 2024).
Sparse attention mechanisms aim to address the aforementioned challenges and reduce computational overhead by approximating dense attention outputs with a subset of key–query interactions (Fu, 2024). Though such methods promise to accelerate long sequence modelling and alleviate memory load while maintaining dense model performance (Jiang et al., 2024), their viability and robustness remain unclear due to a lack of comprehensive large-scale evaluation of state-of-the-art methods. While Li et al. (2025b), Liu et al. (2025b), and Yuan et al. (2024) provide a preliminary exploration of this question, they are limited to a narrow range of configurations (model sizes, sequence lengths, and sparsity levels), specific use cases (such as multi-turn decoding), and rely on datasets with variable sequence lengths that prevent a systematic analysis of length-dependent effects.
@@
In this work, we conduct the largest-scale empirical analysis of training-free [^2] We deliberately concentrate on these approaches because training-based alternatives require either training from scratch (Yuan et al., 2025) or fine-tuning (Nawrot et al., 2024)—both prohibitively expensive in terms of computational resources and requiring access to often confidential training data mixtures. sparse attention methods to date, covering models between 7B and 72B parameters, sequences between 16K to 128K tokens, and sparsity between 0% and 95%. To enable a controlled analysis, we first survey existing approaches, addressing the challenge of comparing rapidly evolving methods whose implementation details often obscure their core design principles. We distil these approaches into four key axes: units of sparsification (blocks/pages or verticals and slashes), importance estimation (fixed or context-aware), budget allocation across layers (uniform or adaptive), and KV cache management (eviction or full cache). Based on this taxonomy, we select six representative patterns spanning these design dimensions and harmonise their implementations, allowing us to rigorously evaluate the distinct effects of each fundamental principle.
@@
For our evaluation, we curate a benchmark suite of 9 long-context tasks designed to systematically probe the influence of key factors on sparse attention performance. These factors include diverse task types (ranging from retrieval to multi-hop tracking and information aggregation), varying naturalness of sequences (synthetic or natural language), and precisely controlled sequence lengths. The importance of these dimensions is underscored by prior work indicating their significant impact on sparse attention effectiveness (Chen et al., 2024; Liu et al., 2024). Alongside established benchmarks (Rajpurkar et al., 2018; Pang et al., 2022; Tseng et al., 2016), we introduce novel, more challenging tasks based on natural language story templates. These complement synthetic benchmarks such as RULER (Hsieh et al., 2024), which measures core skills but whose artificial nature may fail to extrapolate performance on natural data. Our tasks address this gap by instantiating these core skills within a more realistic—and yet fully controllable—natural language setting. All this gives us the opportunity to address fundamental questions that currently remain unresolved:
RQ1: Given a fixed computing budget, should one choose a small dense model, or a large model with sparse attention? We conduct an isoFLOPS analysis finding that the answer depends on sequence length. For short sequences, any increase in either density or size enhances performance; for long sequences, only highly sparse configurations lie on the accuracy—FLOPS Pareto frontier.
RQ2: How far can we push sparsity while statistically guaranteeing that dense attention performance is fully preserved? To assess this, we develop a statistical testing framework reminiscent of distribution-free risk control (Angelopoulos et al., 2022). We find that in general, higher sparsity levels can be achieved during decoding and for larger model sizes. However, this aggregate perspective obscures a crucial caveat: even moderate sparsity levels frequently result in significant performance degradation on at least one task for most configurations.
RQ3: Are there one-size-fits-all methods that are consistently better across a diverse array of long-sequence tasks? We address this question separately for each of the inference phases (prefilling and decoding) and different categories of tasks. We find that no single sparse attention method uniformly excels across all diverse long-sequence tasks, with the ideal unit of sparsification and the choice of whether the budget should be adaptive being task- and phase-specific.
RQ4: Can we establish scaling laws (Brown et al., 2020; Kaplan et al., 2020; Hoffmann et al., 2022) for sparse attention, which can generalise beyond the range of configurations we consider? When holding out model sizes, sequence lengths, or sparsity levels, we can reliably predict their performance through a simple log-linear model, lending credence to the generality of our results.
Overall, our findings support the adoption of sparse attention for enhancing LLM capabilities for processing longer sequences, while emphasising that it requires careful evaluation of trade-offs, particularly for performance-sensitive applications. We release our code at https://github.com/PiotrNawrot/sparse-frontier.
Figure 1: Overview of sparse attention methods for prefilling (left) and generation (right). These methods differ in the units of sparsification (blocks or pages vs. verticals and slashes), importance estimation, and KV cache management strategies. Colours represent query–key interactions preserved at different sparsity levels, while white areas indicate interactions that are not computed.
@@ 2 Training-Free Sparse Attention
Given an input sequence of tokens $X\in\mathbb{R}^{n\times d_{\text{model}}}$ with sequence length $n$ and embedding dimension $d_{\text{model}}$, attention first projects the input into query, key, and value representations: $Q=XW^{Q},K=XW^{K},V=XW^{V}$, where $W^{Q},W^{K},W^{V}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{head}}}$ are the projection matrices, and $d_{\text{head}}$ is the attention head dimensionality. For readability, we omit the multiple attention heads.
Attention computes the output for the $i$-th token as a weighted sum over each value, with weights representing query–key (QK) interactions defined by the attention matrix $A\in\mathbb{R}^{n\times n}$:
$A_{i}=\text{softmax}\left(\frac{Q_{i}K^{\top}}{\sqrt{d_{\text{head}}}}\right)$ (1)
The output for each position is computed as $O_{i}=\sum_{j=1}^{n}A_{ij}V_{j}$, resulting in $O\in\mathbb{R}^{n\times d_{\text{head}}}$.
Text generation with Transformer-based models operates in two distinct phases, each with different computational characteristics affecting attention mechanisms:
Prefilling processes the entire input prompt at once, computing hidden representations for all tokens in parallel. This incurs quadratic computational complexity with respect to sequence length due to the computation of the full attention matrix $A$ in Equation 1.
Decoding generates output tokens auto-regressively, one token at a time. Attention complexity is linear in sequence length per generation step because there is only a single query, but runtime is dominated by high bandwidth memory access for loading key–value (KV) pairs from memory.
By making $A$ sparse, thus computing only a subset of query–key token interactions, sparse attention methods can mitigate these computational bottlenecks: reducing computational overhead during prefilling and memory transfer requirements during decoding.
The effectiveness of sparse attention methods in computation reduction and potentially memory savings is quantified using two key metrics: sparsity, defined as the ratio between the number of attention units (e.g., query–key interactions) that are not computed and the total number of units in dense attention for a given configuration, and compression ratio, defined as $\frac{1}{1-\text{sparsity}}$. For example, sparsity $=0$ corresponds to dense attention where all query–key interactions are computed, while sparsity $=0.9$ means that 90% of interactions are skipped, resulting in a 10$\times$ compression ratio.
We condense the broad variety of existing training-free sparse attention approaches into four main dimensions: unit of sparsification, importance estimation, budget allocation, and KV cache management. We exclude token merging methods (Wang et al., 2024; Nawrot et al., 2024), as they leverage token similarity rather than sparsity. Visualisations of example patterns are shown in Figure 1.
@@ 2.1 Unit of Sparsification
Sparse attention methods differ primarily in the structural units of the attention matrix they prune or retain. Common units include local windows (contiguous regions around each query), vertical columns (tokens globally available to all queries), slashes (tokens at fixed offsets from each query), and blocks (fixed-size tiles of the attention matrix, such as 64$\times$64 tokens). Larger structured units such as blocks or windows offer improved computational efficiency via better memory locality, whereas smaller units allow finer-grained, more precise selection of important information.
@@
Vertical–slash patterns represent another essential class of units. Early sparse attention methods like LM-Infinite (Han et al., 2024) and StreamingLLM (Xiao et al., 2024b) utilised local sliding windows supplemented by prefix tokens shared globally, also known as attention sinks. Extending this approach, Tri-shape (Li et al., 2025b) added full attention for suffix tokens, whereas SnapKV (Li et al., 2024b) introduced dynamically chosen vertical columns. MInference (Jiang et al., 2024) built on this by adding diagonal slashes at arbitrary offsets beyond the local window. [^3] Interestingly, to efficiently compute attention along these diagonals, MInference uses 64×64 blocks aligned with these diagonals rather than computing attention for individual query–key pairs.
@@ 2.2 Importance Estimation
Content-aware methods typically estimate the importance of QK units (tokens, blocks, or diagonals) to retain only the top-k most relevant ones, maximising attention score recall. They use lightweight heuristics such as approximated attention scores from highest-magnitude dimensions (SparQ, Quest; Ribar et al., 2024; Tang et al., 2024) or block-wise pooled token representations (Jiang et al., 2024). Some approaches subsample queries (SampleAttention; Zhu et al., 2024), recognising that recent query tokens often provide better indicators of KV unit importance, as in MInference’s Vertical-Slash pattern (Jiang et al., 2024) and SnapKV (Li et al., 2024b). During decoding, aggregated attention scores from a running average (H2O; Zhang et al., 2023) or the latest query (TOVA; Oren et al., 2024) guide the selection or eviction of KV units, again prioritising units likely to receive the highest attention weights. Some methods incorporate complementary heuristics alongside attention scores, such as vector norms of keys (Devoto et al., 2024) or values (VATP; Guo et al., 2024), recognising that a token’s contribution depends on both its attention score and value vector magnitude, as evident in Equation 1.
@@ 2.3 Budget Allocation
The third key dimension in sparse attention design concerns how to distribute the computational budget across different components of the model like layers and heads given a target sparsity level, which involves fundamental trade-offs between uniform simplicity and adaptive expressivity.
Uniform allocation assumes each head attends to the same number of tokens or blocks, as in Block-Sparse attention (Jiang et al., 2024) and SnapKV (Li et al., 2024b). While computationally simpler, this approach ignores the fact that not all layers and heads contribute equally to model accuracy, and their attention distributions often exhibit diverse sparsity characteristics (Zhang et al., 2024).
Adaptive methods such as PyramidKV (Cai et al., 2024) and PyramidInfer (Yang et al., 2024b) observe that attention scores’ entropy decreases with layer depth, suggesting that early layers require larger budgets than deeper ones. Similarly, Mixture of Sparse Attention (MoA; Fu et al., 2024) employs an automatic budget-fitting procedure using first-order Taylor approximations to optimally distribute the global attention budget across layers. Complementing this inter-layer approach, Ada-KV (Feng et al., 2024) focuses on flexible allocation within each layer by selecting the top-$(k\times h)$ tokens per layer (where $h$ is the number of heads), redistributing the total cache size so that more critical heads retain additional keys while less critical ones are aggressively pruned.
Adaptive threshold-based allocation represents the most flexible approach, relaxing the constraint of a fixed global budget altogether. Methods like Twilight (Lin et al., 2025), FlexPrefill (Lai et al., 2025), Tactic (Zhu et al., 2025), and SampleAttention (Zhu et al., 2024) establish coverage thresholds (e.g., capturing 95% of the attention weight mass) and allow each head to dynamically select as many units as necessary to reach these thresholds. Consequently, heads exhibiting diffuse (high-entropy) attention naturally consume a larger share of the budget, whereas heads with sharply peaked attention consume fewer tokens. These methods also incorporate a minimum budget parameter to make them robust to edge cases of extreme sparsity (Chen et al., 2024).
@@ 2.4 KV Cache Management
The final dimension distinguishing sparse attention methods concerns their approach to KV cache management, particularly critical during the memory-bound decoding phase. While prefilling benefits primarily from computational speed-ups, decoding performance is often limited by the memory required to store and access the KV cache of past tokens. Sparse attention methods for decoding address this bottleneck through two main strategies, creating a fundamental trade-off between memory footprint reduction and information fidelity.
One strategy involves KV cache eviction, where methods like H2O (Zhang et al., 2023) or SnapKV (Li et al., 2024b) permanently discard selected tokens or blocks from the cache based on estimated importance. This directly reduces the memory footprint and data transfer costs, enabling longer sequences within fixed memory budgets and potentially improving hardware utilisation. However, this approach sacrifices information fidelity, as discarded tokens cannot be recovered if they become relevant later in the generation process. Robustly identifying tokens for eviction is challenging, as the set of important tokens can vary significantly across generation steps, particularly with shifting contexts (Li et al., 2025b).
Alternatively, other methods maintain the full KV cache but optimise computation by selectively loading only necessary KV pairs from memory during attention calculation, exemplified by Quest (Tang et al., 2024) and SparQ (Ribar et al., 2024). While they may incur a small memory overhead for auxiliary data structures used for importance estimation (e.g., page summaries), they avoid the information loss inherent in eviction. This often allows for effectively operating at higher sparsity levels compared to eviction-based methods (Li et al., 2025b), although they do not reduce the peak memory requirement for storing the cache itself.

### v2

@@ Abstract
Sparse attention offers a promising strategy to extend long-context capabilities in Transformer LLMs, yet its efficiency–accuracy trade-offs remain unclear due to the lack of comprehensive evaluation. We address this gap with the largest-scale empirical analysis to date of training-free sparse attention, evaluating six methods across multiple model families and sizes, sequences up to 128K tokens, and sparsity levels up to 0.95 (i.e., $1/20$ attention budget) on nine diverse tasks. We first organise the rapidly evolving landscape of sparse attention methods into a taxonomy along four design axes. Our analysis then yields actionable insights: 1) sparse attention is effective—larger sparse models outperform smaller dense ones at equivalent cost, improving the Pareto frontier; 2) due to computational constraints, token-to-page importance estimation is unfeasible during prefilling, where the choice of an alternative solution (global-to-token or block-to-block) depends on the task, but is possible during decoding, enabling better generalisation and tolerance to higher sparsity; 3) longer sequences tolerate higher sparsity, suggesting that fixed-budget methods in production are suboptimal. Together, these findings provide practical guidance for deploying sparse attention and methodological recommendations for future evaluations. Our code is available at https://github.com/PiotrNawrot/sparse-frontier.
@@ 1 Introduction
The ability to model long sequences in large language models (LLMs) lies at the heart of long-context processing (Liu et al., 2025a) and inference-time scaling (Snell et al., 2024; Muennighoff et al., 2025). The fundamental bottleneck for this ability is the self-attention mechanism (Bahdanau et al., 2015; Vaswani et al., 2017): during the prefilling stage, its computational complexity scales quadratically with sequence length—hence, ballooning time-to-first-token and deployment cost (Jiang et al., 2024). In the decoding phase, the key–value (KV) cache grows linearly with sequence length, and the need to load from memory this expanding cache for each generation step dominates the runtime (Nawrot et al., 2024).
Sparse attention mechanisms aim to address these challenges by approximating dense attention outputs with only a subset of query–key interactions (Fu, 2024). These span both training-based variants—such as DMS (Łańcucki et al., 2025), DeepSeek’s NSA (Yuan et al., 2025), and SWA used in OpenAI’s gpt-oss and Google’s Gemma 3—and training-free methods that operate directly on pretrained models, such as Vertical-Slash (Jiang et al., 2024) deployed in Qwen 2.5-1M (Yang et al., 2025) and integrated into vLLM. Sparse attention has not only seen widespread adoption in industry, but also in the research community: over 150 papers with “sparse attention” in the title were submitted to arXiv between January 2025 and January 2026. Despite this popularity, the viability and robustness of sparse attention remain unclear due to a lack of comprehensive large-scale evaluation.
@@
In this work, we carry out the largest-scale empirical analysis to date of training-free sparse attention methods, covering three model families (Qwen 2.5, Llama 3.1, Gemma 3) with sizes between 4B and 72B parameters, sequences between 16K to 128K tokens, and sparsity levels up to 0.95 (i.e., $1/20$ attention budget). To enable a controlled analysis, we first survey existing approaches, addressing the challenge of comparing rapidly evolving methods whose implementation details often obscure their core design principles. We distil these approaches into four key axes: units of sparsification (blocks/pages or verticals and slashes), importance estimation (fixed or context-aware), budget allocation across layers (uniform or adaptive), and KV cache management (eviction or full cache). Based on this taxonomy, we select six representative methods spanning these design dimensions and harmonise their implementations, allowing us to rigorously evaluate their distinct effects.
We focus specifically on training-free sparse attention because training-based alternatives require prohibitive computational resources and possibly access to proprietary training data (Nawrot et al., 2024). While this is a limitation, we expect insights from our training-free analysis to transfer to training-based methods given their similarity.***For instance, Quest (Tang et al., 2024) and NSA (Yuan et al., 2025) both use page-based selection for sparse decoding. While previously Li et al. (2025), Liu et al. (2025b), and Yuan et al. (2024) provided a preliminary exploration of training-free methods, they covered limited configurations, specific use cases, or did not control for sequence lengths, hindering a systematic analysis (see Appendix C for a detailed comparison).
@@
For our evaluation, we curate a benchmark suite of 9 long-context tasks designed to systematically probe the influence of key factors on sparse attention performance. These factors include diverse task types (ranging from retrieval to multi-hop variable tracking and information aggregation), varying naturalness of sequences (synthetic or natural language), and precisely controlled sequence lengths. The importance of these dimensions is underscored by prior work indicating their significant impact on sparse attention effectiveness (Chen et al., 2024; Liu et al., 2024). Alongside established benchmarks (Rajpurkar et al., 2018; Pang et al., 2022; Tseng et al., 2016), we introduce novel, more challenging tasks based on natural language story templates. These complement synthetic benchmarks like RULER (Hsieh et al., 2024), whose results may fail to extrapolate to realistic data, by evaluating core skills in a controllable yet realistic natural language setting. All this provides us with a toolbox to address fundamental questions that currently remain unresolved:
RQ1: Is sparse attention effective? (Section 4.1) An isoCost analysis reveals that sparsification enables larger sparse models to outperform smaller dense ones at equivalent cost (i.e., FLOPs during prefilling and memory reads during decoding). For long sequences, only high-sparsity configurations lie on the Pareto frontier.
RQ2: Which sparse attention method should practitioners use? (Section 4.2) Prefill and decoding phases display different trends. During prefilling, computational constraints force a choice between fine-grained token selection (Vertical-Slash) and block-based selection (Block-Sparse)—neither of which generalises across all tasks. During decoding, token-to-page selection (e.g., Quest) is computationally feasible, enabling greater flexibility and higher compression tolerance than prefilling.
RQ3: How does the sequence length affect tolerance to sparse attention? (Section 4.3) Longer sequences permit higher sparsity while maintaining accuracy, consistently across model families. This suggests that fixed-budget methods deployed in production are suboptimal, and future designs should adapt sparsity to sequence length.
Overall, our findings provide practical guidance for deploying sparse attention and methodological recommendations for future evaluations in this rapidly evolving field.
@@ 2 Training-Free Sparse Attention
The self-attention mechanism computes query $Q$, key $K$, and value $V$ representations from an input sequence $X\in\mathbb{R}^{n\times d}$. The output $O_{i}$ for the $i$-th token is a weighted sum of values, $O_{i}=\sum_{j=1}^{n}A_{ij}V_{j}$, where the attention weights $A_{ij}$ are derived from scaled dot-products between queries and keys: $A_{i}=\text{softmax}(Q_{i}K^{\top}/\sqrt{d})$. We omit multi-head details for brevity.
Transformer-based text generation involves two phases. Prefilling processes the entire input sequence, computing the lower-triangular part of the $n\times n$ attention matrix $A$, leading to $O(n^{2})$ complexity. Decoding generates tokens auto-regressively. While attention is $O(n)$ per step (single query), loading the expanding Key-Value (KV) cache from memory becomes the main bottleneck.
Sparse attention methods reduce these costs by computing only a subset of QK interactions, making $A$ sparse. This lowers computational load during prefilling and memory transfers during decoding. We quantify the effectiveness of these methods using sparsity—the fraction of non-computed QK interactions. Equivalently, sparsity of 11/k1-1/k corresponds to retaining only a $1/k$ fraction of the attention interactions. For instance, sparsity 0.9 (or equivalently, $1/10$ attention budget) means computing only 10% of the original QK interactions.
The speedup from sparse attention depends on how much of total cost is attention. Since attention scales quadratically while other components (MLP, embeddings) scale linearly with sequence length, attention dominates at longer contexts—yielding greater benefits from sparsification. Models with built-in architectural sparsity, such as sliding-window or linear attention layers, have lower baseline attention ratios and require longer sequences for additional sparsification to provide comparable gains (Appendix B).
We categorise training-free sparse attention methods along four axes: unit of sparsification, importance estimation, budget allocation, and KV cache management. We exclude token merging methods (Wang et al., 2024; Nawrot et al., 2024), which do not rely on sparsity.
@@ 2.1 Unit of Sparsification
Sparse attention methods differ primarily in the structural units of the attention matrix they prune or retain. Common units include local windows (contiguous regions around each query), vertical columns (tokens globally available to all queries), slashes (tokens at fixed offsets from each query), and blocks (fixed-size tiles of the attention matrix, such as $\times64$ tokens). Larger structured units such as blocks or windows offer improved computational efficiency via better memory locality, whereas smaller units allow finer-grained, more precise selection of important information.
@@
Vertical–slash patterns represent another essential class of units. Early sparse attention methods like LM-Infinite (Han et al., 2024) and StreamingLLM (Xiao et al., 2024b) utilised local sliding windows supplemented by prefix tokens shared globally, also known as attention sinks. Extending this approach, Tri-shape (Li et al., 2025) added full attention for suffix tokens, whereas SnapKV (Li et al., 2024b) introduced dynamically chosen vertical columns. MInference (Jiang et al., 2024) built on this by adding diagonal slashes at arbitrary offsets beyond the local window.Interestingly, to efficiently compute attention along these diagonals, MInference uses 64×64 blocks aligned with these diagonals rather than computing attention for individual query–key pairs. For a visual illustration of these attention patterns, see Figure 4 in the Appendix.
@@ 2.2 Importance Estimation
Content-aware methods typically estimate the importance of QK units (tokens, blocks, or diagonals) to retain only the top-k most relevant ones, maximising attention score recall. They use lightweight heuristics such as approximated attention scores from highest-magnitude dimensions (SparQ; Ribar et al., 2024) or block-wise pooled token representations (Jiang et al., 2024). Some approaches subsample queries (SampleAttention; Zhu et al., 2024), recognising that recent query tokens often provide better indicators of KV unit importance, as in MInference’s Vertical-Slash (Jiang et al., 2024) and SnapKV (Li et al., 2024b). During decoding, aggregated attention scores (H2O; Zhang et al., 2023) or the latest query (TOVA; Oren et al., 2024) guide the selection of KV units, again prioritising units likely to receive high attention weights. Some methods incorporate complementary heuristics alongside attention scores, such as norms of keys (Devoto et al., 2024) or values (Guo et al., 2024).
Critically, the cost of sparse attention includes both the sparse operation and importance estimation overhead. During prefilling, per-cell importance estimation would require quadratic cost, so methods must either select fine-grained units globally (e.g., vertical columns shared across all queries) or use coarser block-to-block selection. More flexible unit selection improves accuracy but increases estimation cost. During decoding, per-query selection is feasible since only one query is processed per step, enabling methods like Quest to perform finer token-to-block selection and tolerate higher compression (Section 4.2).
@@ 2.3 Budget Allocation
The third dimension in sparse attention design is budget allocation: distributing computational resources across model components (layers and heads) for a target sparsity. This involves a trade-off between uniform simplicity and adaptive expressivity.
Uniform allocation assigns an equal budget (tokens or blocks) to each head as in Block-Sparse (Jiang et al., 2024) and SnapKV (Li et al., 2024b). This is computationally simple but overlooks that layers and heads contribute differently to accuracy and have diverse attention sparsity (Zhang et al., 2024).
Adaptive methods vary budget allocation. PyramidKV (Cai et al., 2024) and PyramidInfer (Yang et al., 2024b) observe that attention score entropy decreases with layer depth, allocating larger budgets to early layers. Mixture of Sparse Attention (MoA; Fu et al., 2024) uses Taylor approximations to optimally distribute the global budget across layers. Within layers, Ada-KV (Feng et al., 2024) flexibly allocates by selecting top-$(k\times h)$ tokens (where $h$ is head count), allowing critical heads to retain more keys while pruning others. Threshold-based allocation offers maximum flexibility by removing a fixed global budget. Methods like Twilight (Lin et al., 2025), FlexPrefill (Lai et al., 2025), Tactic (Zhu et al., 2025), and SampleAttention (Zhu et al., 2024) set coverage thresholds (e.g., 95% of attention mass). Each head dynamically selects units to meet these thresholds, allowing high-entropy attention heads to consume more budget and the overall budget to vary per sample.
@@ 2.4 KV Cache Management
The final dimension distinguishes methods based on KV cache management during decoding.
KV cache eviction methods (e.g., H2O (Zhang et al., 2023), SnapKV (Li et al., 2024b)) permanently discard selected tokens based on estimated importance, directly reducing memory footprint but sacrificing information fidelity as discarded tokens cannot be recovered.
Full KV cache retention methods (e.g., Quest (Tang et al., 2024), SparQ (Ribar et al., 2024)) maintain the entire cache but optimize computation by selectively loading only necessary KV pairs during attention calculation. While incurring small memory overhead for auxiliary data structures needed for importance estimation, they avoid information loss and can operate effectively at higher sparsity levels compared to eviction-based methods, though they do not reduce peak memory requirements.